OpenStudy (anonymous):

Give the function f(x)= 3x+3, find the value of one third. A. 3 B. 1/3 C. 4 D. 1

hero (hero):

You mean, find f(1/3)

mathslover (mathslover):

did u mean to find f(1/3) ?

OpenStudy (anonymous):

c

OpenStudy (lgbasallote):

i can randomly say A...doesnt make me any more wrong than @helpmexoxo

mathslover (mathslover):

plug in 1/3 in the place of x what u get?

hero (hero):

f(x) = 3x + 3 f(1/3) = 3(1/3) + 3 f(1/3) = 3/3 + 3 f(1/3) = 1 + 3 f(1/3) = 4 C

hero (hero):

Gulp, I thought @helpmexoxo was the original poster.